OpenStudy (anonymous):

Convert the given coordinates to the indicated ordered pair: (1, - sqrt 3) to (r, theta). I have found r to be 2 but I am having trouble finding theta. Please help!!!

OpenStudy (freckles):

 so just do arctan(y/x) foe theta

OpenStudy (anonymous):

Okay, thank you so much (again). Also, did I get r correct?

OpenStudy (anonymous):

Also, when I do arctan(y/x) I get -1.047... should I have a more accurate answer in terms of pi for theta?

OpenStudy (anonymous):

Okay! I think I've found the answer.... the converted pair would be \[(2, -\pi/3)\] ?

OpenStudy (freckles):

sorry notifications isn't working properly

OpenStudy (freckles):

r^2=1+3=4 r=2 so good and yes theta=-pi/3 now if there is no restriction on r or theta then you are done

OpenStudy (anonymous):

Well actually, is there any restriction since it must be in the fourth quadrant?

OpenStudy (freckles):

your r and theta is fine

OpenStudy (freckles):

Sometimes the teacher wants a positive theta

OpenStudy (freckles):

or a negative r

OpenStudy (freckles):

but if there is nothing like that given then you are done

OpenStudy (anonymous):

yeah, it just says to convert. Thank you so much!!

OpenStudy (freckles):

np